# Semantic Prior Distillation with Vision Foundation Model for Enhanced Rapid Bone Scintigraphy Image Restoration

### v2

@@ Abstract
Rapid bone scintigraphy is an essential tool for diagnosing skeletal diseases and tumor metastasis in pediatric patients, as it reduces scan time and minimizes patient discomfort. However, rapid scans often result in poor image quality, potentially affecting diagnosis due to reduced resolution and detail, which make it challenging to identify and evaluate finer anatomical structures. To address this issue, we propose the first application of SAM-based semantic priors for medical image restoration, leveraging the Segment Anything Model (SAM) to enhance rapid bone scintigraphy images in pediatric populations. Our method comprises two cascaded networks, $f^{IR1}$ and $f^{IR2}$, augmented by three key modules: a Semantic Prior Integration (SPI) module, a Semantic Knowledge Distillation (SKD) module, and a Semantic Consistency Module (SCM). The SPI and SKD modules incorporate domain-specific semantic information from a fine-tuned SAM, while the SCM maintains consistent semantic feature representation throughout the cascaded networks. In addition, we will release a novel Rapid Bone Scintigraphy dataset called RBS, the first dataset dedicated to rapid bone scintigraphy image restoration in pediatric patients. RBS consists of 137 pediatric patients aged between 0.5 and 16 years who underwent both standard and rapid bone scans. The dataset includes scans performed at 20 cm/min (standard) and 40 cm/min (rapid), representing a 2× acceleration. We conducted extensive experiments on both the publicly available endoscopic dataset and RBS. The results demonstrate that our method outperforms all existing methods across various metrics, including PSNR, SSIM, FID, and LPIPS.
@@ 3.5 Semantic Consistency Module
To ensure consistent semantic feature representation across both IR models, we incorporate the Semantic Consistency Module. This module plays a crucial role in maintaining the semantic coherence of the restored images, which is essential for high-quality image restoration. The SCM utilizes a pre-trained VGG network to extract semantic feature maps from both $I_{HQ}^{1}$ and $I_{HQ}^{2}$, ensuring semantic relationship preservation and alignment throughout the restoration process.
The SCM operates on the principle that semantically similar regions in the input image should maintain their relationships in the restored output, regardless of low-level pixel variations. To achieve this, we first extract high-level features from the restored images using a pre-trained VGG network:
$F_{VGG}^{1}=VGG(I_{HQ}^{1}),\quad F_{VGG}^{2}=VGG(I_{HQ}^{2}),$ (8)
where $F_{VGG}^{1}$ and $F_{VGG}^{2}$ are the feature maps extracted from $I_{HQ}^{1}$ and $I_{HQ}^{2}$ respectively.
To focus on specific semantic regions, we utilize the semantic mask $M$ obtained from SAM. We apply this mask to the VGG features to obtain region-specific semantic representations:
$F_{VGG}^{1}(n)=F_{VGG}^{1}\odot M_{n},\quad F_{VGG}^{2}(n)=F_{VGG}^{2}\odot M_{n},$ (9)
where $M_{n}$ is the binary mask for the $n$-th semantic region, and $\odot$ denotes element-wise multiplication.
Next, we compute the semantic relationships between different regions. This is done by calculating the normalized dot product between the feature vectors of different regions:
$R_{VGG}^{1}(n_{1},n_{2})=\frac{F_{VGG}^{1}(n_{1})\cdot F_{VGG}^{1}(n_{2})}{\|F_{VGG}^{1}(n_{1})\|_{2}\|F_{VGG}^{1}(n_{2})\|_{2}},$ (10)
where $\cdot$ denotes the dot product and $\|\cdot\|_{2}$ is the L2 norm. $R_{VGG}^{2}(n_{1},n_{2})$ is computed similarly for $F_{VGG}^{2}$.
The SCM loss is then defined as the difference between these semantic relationship matrices:
$\mathcal{L}_{SCM}=\sum_{n_{1},n_{2}=1,n_{1}\neq n_{2}}^{N}\frac{|R_{VGG}^{1}(n_{1},n_{2})-R_{VGG}^{2}(n_{1},n_{2})|}{N^{2}-N},$ (11)
where $N$ is the total number of semantic regions. This loss encourages the outputs of both IR models to maintain consistent semantic relationships, even if they differ in low-level details. By integrating the SCM into our framework, we achieve a balance between pixel-level accuracy and semantic fidelity in image restoration. This leads to restored images that not only have improved visual quality but also maintain the semantic integrity of the original scene, which is essential for many downstream computer vision tasks and human perception.

### v3

# Rapid Bone Scintigraphy Enhancement via Semantic Prior Distillation from Segment Anything Model

@@ Abstract
Rapid bone scintigraphy is crucial for diagnosing skeletal disorders and detecting tumor metastases in children, as it shortens scan duration and reduces discomfort. However, accelerated acquisition often degrades image quality, impairing the visibility of fine anatomical details and potentially compromising diagnosis. To overcome this limitation, we introduce the first application of SAM-based semantic priors for medical image restoration, utilizing the Segment Anything Model (SAM) to enhance pediatric rapid bone scintigraphy. Our approach employs two cascaded networks, $f^{IR1}$ and $f^{IR2}$, supported by three specialized modules: a Semantic Prior Integration (SPI) module, a Semantic Knowledge Distillation (SKD) module, and a Semantic Consistency Module (SCM). The SPI and SKD modules inject domain-specific semantic cues from a fine-tuned SAM, while the SCM preserves coherent semantic feature representations across both cascaded stages. Moreover, we present RBS, a novel Rapid Bone Scintigraphy dataset comprising paired standard (20 cm/min) and rapid (40 cm/min) scans from 137 pediatric patients aged 0.5–16 years, making it the first dataset tailored for pediatric rapid bone scintigraphy restoration. Extensive experiments on both a public endoscopic dataset and our RBS dataset demonstrate that our method consistently surpasses existing techniques in PSNR, SSIM, FID, and LPIPS metrics.
@@ 3.5 Semantic Consistency Module
To enforce consistent semantic feature representation between the two restoration branches, we introduce a Semantic Consistency Module (SCM). Let the high-resolution outputs from the two restoration paths be $X^{A}$ and $X^{B}$. We first extract deep feature maps using a fixed, pretrained convolutional network $\phi$ (VGG):
$H^{A}=\phi\bigl{(}X^{A}\bigr{)},\quad H^{B}=\phi\bigl{(}X^{B}\bigr{)},$ (8)
where $H^{A},H^{B}\in\mathbb{R}^{C\times H\times W}$ are feature tensors with $C$ channels and spatial dimensions $H\times W$.
Let $\{S_{k}\}_{k=1}^{K}$ be $K$ binary masks produced by a segmentation model (SAM), where each mask $S_{k}\in\{0,1\}^{H\times W}$ isolates one semantic region. We apply these masks to the feature maps and then vectorize:
$h^{A}_{k}=\mathrm{vec}\bigl{(}H^{A}\odot S_{k}\bigr{)},\quad h^{B}_{k}=\mathrm{vec}\bigl{(}H^{B}\odot S_{k}\bigr{)},$ (9)
where $\odot$ denotes element-wise multiplication and $\mathrm{vec}(\cdot)$ flattens a tensor into a vector.
Next, for every pair of regions $(i,j)$, we compute the cosine similarity between their feature vectors to capture high-level semantic relationships:
$c^{A}_{ij}=\frac{\langle h^{A}_{i},\,h^{A}_{j}\rangle}{\|h^{A}_{i}\|_{2}\,\|h^{A}_{j}\|_{2}},\quad c^{B}_{ij}=\frac{\langle h^{B}_{i},\,h^{B}_{j}\rangle}{\|h^{B}_{i}\|_{2}\,\|h^{B}_{j}\|_{2}}.$ (10)
Finally, we define the semantic alignment loss to encourage the two branches to preserve the same relational structure among regions:
$\mathcal{L}_{\mathrm{SCM}}=\frac{1}{K(K-1)}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{K}\bigl{|}\,c^{A}_{ij}-c^{B}_{ij}\bigr{|}.$ (11)
By integrating $\mathcal{L}_{\mathrm{SCM}}$ into the overall training objective, the model not only achieves pixel-level accuracy but also maintains semantic consistency across restored images, leading to higher visual fidelity and improved performance on downstream tasks.